Manuel has a boat that can move at a speed of 15 km/h in still water. He rides 140 km downstream in a river in the same time it takes to ride 35 km upstream. What is the speed of the river?
Answer by mananth(16946)

Boat speed = 15 km/h
current speed = x km/h

speed against current 15 - x
speed with current 15 + x

Distance against current 35 miles
Distance with current 140 miles

t=d/r time against = time with current -
35 / ( 15 - x )= 140 / ( 15 + x )
35 ( 15 + x )= 140 ( 15 -x )
525 + 35 x = 2100 - -140 x
140 x 35 x 2100 + -525
175 x = 1575
/ 175
x = 9 km/h current speed